If three coins are tossed, what are the mean, variance, and standard deviation of the number of heads that occur?


1
Expert's answer
2022-03-18T10:26:31-0400

The number of heads is "X\sim Bin(3,0.5)"

Hence

"EX=3\cdot 0.5=1.5\\
DX=3\cdot 0.5 \cdot (1-0.5)=0.75\\
\sigma X=\sqrt{DX}=0.866"